Jefferson Construction builds custom houses and uses a job costing system with two direct cost categories (direct materials and direct labor). Overhead is in one pool, allocated by direct labor dollars. Estimated overhead for the current year is $9,900,000; expected direct labor hours are 550,000; expected direct material dollars are $10,000,000; expected direct labor dollars are $6,600,000.

Jefferson is interested in a comparison between two jobs (where two different styles of home were built) because resources will be limited next year, and the company will have to start turning down less profitable contracts.

If one Madison and one Monroe home above were the only projects done during the current month and actual overhead for the current month was $250,000, was overhead under- or over-applied? By how much?

Madison Monroe
DM 90,000 104,000
DL 85,000 75,000
Direct Labor Hours 5,450 5,000
Answer #1
Predetermined overhead rate = Estimated overhead / Estimated direct labor dollars = 9900000 / 6600000 = 1.5 per direct labor dollar
Actual Total direct labor dollars = 85000 + 75000 = 160000
Overhead applied = Predetermined overhead rate * Actual Total direct labor dollars = 1.5 * 160000 = 240000
Actual overhead incurred 250000
As the overhead applied is less than the actual overhead incurred, the overhead is under-applied
Underapplied overhead = Actual overhead incurred - Overhead applied = 250000 - 240000 = 10000